OpenStudy (anonymous):

If m is directly proportional to r^3, and m=1/12 when r=1/2, what is the value of m when r=3/2?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Directrix

OpenStudy (zzr0ck3r):

mk=r^3 for some number k let m = 1/12 and r = 1/2 k/12 = 1/8 k = 12/8 = 3/4 so (3/4)m=r^3 let r = (3/2) (3/4)m = (3/2)^3 = 27/8 (3/4)m = 27/8 m = (4/3)(27/8) = 4(9/8) = 9/2

Directrix (directrix):

@zzr0ck3r Why would the initial equation not be the following: m=kr^3 for some number k I'm confused on that.

OpenStudy (zzr0ck3r):

well its just a number so m=kr^3 would just be m(1/k) = r^3 just l = 1/k and we have ml=r^3

OpenStudy (zzr0ck3r):

does direct proportion mean integer?

Directrix (directrix):

What about this: m* (1/k) =r^3 for some number k Would that be correct?

OpenStudy (zzr0ck3r):

I think, but k is just a number so no reason to right 1/k. just call that k.

OpenStudy (zzr0ck3r):

write*

OpenStudy (anonymous):

Neither of your answers are on the answer choices

OpenStudy (anonymous):

Plus, I don't know the answer

Directrix (directrix):

@sakigirl That's what mathematics is all about - getting answers not the first time around but sometime around. Cogitation is everything. :)

OpenStudy (anonymous):

Haha, that's very true

OpenStudy (zzr0ck3r):

woops mk=r^3 for some number k let m = 1/12 and r = 1/2 k/12 = 1/8 k = 12/8 = 3/2 NOT (3/4) so m(3/2) = r^3 let r = (3/2) m(3/2) = 27/8 m = (2/3)(27/8) = 2(9/8) = 9/4 is 9/4 a choice?

OpenStudy (anonymous):

Yes!

OpenStudy (zzr0ck3r):

if not try the other way lol m = k * r^3 let m = 1/12 and r = 1/2 1/12 = k * 1/8 k = 2/3 r = 3/2 m = (2/3)(27/8) = 9/4 so it does not matter if we do m = kr^3 or mk=r^3 @Directrix I guess that makes sense....I guess lol

Directrix (directrix):

Well, I started out this way: r^3 * 1/12 = m * (3/2)^3 (1/2)^3 * (1/12) = m * (3/2)^3 1/8 * 1/12 = m * 27/8 1/8 * 1/12 * 8/27 = m At that point, m went off track. I did not account for the constant of proportionality.

OpenStudy (zzr0ck3r):

ahh